# Improving Video Generation with Human Feedback

### v1

@@ 3.2 Reward Model Learning
where $\theta$ is a parameter that controls the tendency toward ties, with a larger $\theta$ indicating a higher probability of ties. We empirically set $\theta=5.0$ and optimize the BTT model by minimizing the negative log-likelihood (NLL):
$\mathcal{L}_{\text{BTT}}=-\mathbb{E}_{({\bm{y}},{\bm{x}}_{0}^{A},{\bm{x}}_{0}^{B})\sim\mathcal{D}}\Big{[}$ $\mathbbm{1}({\bm{x}}_{0}^{A}\succ{\bm{x}}_{0}^{B})\log P({\bm{x}}_{0}^{A}\succ{\bm{x}}_{0}^{B}|{\bm{y}})+$ (6)
$\mathbbm{1}({\bm{x}}_{0}^{B}\succ{\bm{x}}_{0}^{A})\log P({\bm{x}}_{0}^{B}\succ{\bm{x}}_{0}^{A}|{\bm{y}})+$
$\mathbbm{1}({\bm{x}}_{0}^{A}={\bm{x}}_{0}^{B})\log P({\bm{x}}_{0}^{A}={\bm{x}}_{0}^{B}|{\bm{y}})\Big{]}$
We train both the Bradley-Terry reward model and the Bradley-Terry-With-Tied reward model under the same setting and visualize the distribution of $\Delta r=r({\bm{x}}_{0}^{A},{\bm{y}})-r({\bm{x}}_{0}^{B},{\bm{y}})$ on the validation set using the best-performing checkpoints. As shown in the boxplot in Fig. 4, while the BT model effectively manages chosen/rejected pairs, it struggles to accurately differentiate tied pair. A large portion of ties exhibit a larger $\Delta r$ than certain chosen/rejected pairs, underscoring the BT model’s limitations in handling ambiguous cases. In contrast, the BTT model learns a more generalized decision boundary, enhancing it to distinguish tied pairs from clear chosen/rejected pairs based on $\Delta r$. This improvement stems from the explicit modeling of ties in Eq. 3, which allows the model to incorporate this often-overlooked category, resulting in more robust performance.
Figure 4: Visualization of the $\Delta r$ distribution for the BT reward model (Left) and the BTT reward model (Right). The BTT model effectively distinguishes tie pairs from chosen/rejected pairs.
Separation for Decoupling.
When adopting LLMs/MLLMs for multi-attribute reward modeling, a common practice (Ouyang et al., 2022; Touvron et al., 2023; He et al., 2024) involves extracting the last token from the final layer and applying a trainable linear projection to predict scores across multiple dimensions. While this approach enables efficient multi-dimensional scoring, it inherently induces coupling among dimensions, particularly between context-agnostic and context-aware attributes. For example, such a model may inadvertently rely on the input prompt when evaluating visual quality, causing the same video sample to receive different visual quality scores depending on its prompt. Although this coupling may be inherited from the correlations within the annotated data, it compromises the interpretability of reward scores and introduces challenges for the evaluation or optimization of individual dimensions.
To decouple context-agnostic dimensions from contextual influences, we introduce separate special tokens for each dimension. Specially, three unique query tokens are added to represent the scores across the three evaluation dimensions. In the input template of our reward models, the context-agnostic tokens (i.e., VQ, MQ) are placed after the video but before the prompt, while the context-aware special tokens (i.e., TA) are positioned after the prompt. We then extract the feature representations from the final layer corresponding to each special token’s position and apply a shared linear head to predict the associated dimension scores. Owing to the autoregressive model’s causal attention mechanism, the context-agnostic tokens exclusively attend to the video content, whereas the context-aware tokens attend to both the video and the prompt. This strategy effectively separates the focus of each token, enabling a more robust performance of each dimension and independent evaluation for context-agnostic dimensions. We provide a full version of our input prompt template in appendix B.
@@ 4 Video Alignment
Let ${\bm{x}}_{0}\sim q({\bm{x}}_{0})$ denote a real data sample drawn from the true data distribution, and let ${\bm{x}}_{1}\sim p({\bm{x}}_{1})$ denote a noise sample, where ${\bm{x}}_{0},{\bm{x}}_{1}\in\mathbb{R}^{d}$. Recent advanced image-generation models (e.g., (Esser et al., 2024)) and video-generation models (e.g., (Kong et al., 2024)) adopt the Rectified Flow framework (Liu et al., 2022), which defines the “noised” data ${\bm{x}}_{t}$ as
${\bm{x}}_{t}=(1-t)\,{\bm{x}}_{0}\;+\;t\,{\bm{x}}_{1},$
for $t\in[0,1]$. Then we can train a transformer model to directly regress the velocity field ${\bm{v}}_{\theta}({\bm{x}}_{t},t)$ by minimizing the Flow Matching objective (Lipman et al., 2022; Liu et al., 2022):
@@ 4.1 Flow-DPO
Consider a fixed dataset $\mathcal{D}=\{{\bm{y}},{\bm{x}}_{0}^{w},{\bm{x}}_{0}^{l}\}$, where each sample consists of a prompt ${\bm{y}}$ and two videos ${\bm{x}}_{0}^{w}$ and ${\bm{x}}_{0}^{l}$ generated by a reference model $p_{\text{ref}}$, with human annotations indicating that ${\bm{x}}_{0}^{w}$ is preferred over ${\bm{x}}_{0}^{l}$ (i.e., ${\bm{x}}_{0}^{w}\succ{\bm{x}}_{0}^{l}$). The goal of RLHF is to learn a conditional distribution $p_{\theta}({\bm{x}}_{0}\mid{\bm{y}})$ that maximizes a reward model $r({\bm{x}}_{0},{\bm{y}})$, while controlling the regularization term (KL-divergence) from the reference model $p_{\text{ref}}$ via a coefficient $\beta$:
$\max_{p_{\theta}}$ $\mathbb{E}_{{\bm{y}}\sim\mathcal{D}_{c},{\bm{x}}_{0}\sim p_{\theta}({\bm{x}}_{0}\mid{\bm{y}})}\left[r({\bm{x}}_{0},{\bm{y}})\right]$
$-\beta\,\mathbb{D}_{\text{KL}}\left[p_{\theta}({\bm{x}}_{0}\mid{\bm{y}})\,\|\,p_{\text{ref}}({\bm{x}}_{0}\mid{\bm{y}})\right]$ (7)
In this context, $p_{\text{ref}}$ corresponds to $p_{\text{ref},0}$ and $p_{\theta}$ corresponds to $p_{\theta,0}$. For the sake of simplicity, the timestep subscripts are omitted. However, RLHF is usually unstable and resource intensive (Rafailov et al., 2024). To address this issue, Diffusion-DPO aligns diffusion models with human preferences by directly solving the above RLHF objective (Eq. 4.1) analytically. It interprets alignment as a classification problem, and optimizes a policy to satisfy human preferences by supervised training. For simplicity, we omit the conditioning prompt ${\bm{y}}$ in the following equations. The Diffusion-DPO objective $\mathcal{L}_{\text{DD}}(\theta)$ is given by:
$-\mathbb{E}\Bigg{[}\log\sigma\Bigg{(}$ $-\frac{\beta}{2}\Big{(}\|{\bm{\epsilon}}^{w}-{\bm{\epsilon}}_{\theta}({\bm{x}}_{t}^{w},t)\|^{2}-\|{\bm{\epsilon}}^{w}-{\bm{\epsilon}}_{\text{ref}}({\bm{x}}_{t}^{w},t)\|^{2}$
$\quad-\bigl{(}\|{\bm{\epsilon}}^{l}-{\bm{\epsilon}}_{\theta}({\bm{x}}_{t}^{l},t)\|^{2}-\|{\bm{\epsilon}}^{l}-{\bm{\epsilon}}_{\text{ref}}({\bm{x}}_{t}^{l},t)\|^{2}\bigr{)}\Big{)}\Bigg{)}\Bigg{]}$ (8)
where ${\bm{x}}_{t}^{*}=(1-t)\ {\bm{x}}_{0}^{*}+t\ {\bm{\epsilon}}$, ${\bm{\epsilon}}\sim\mathcal{N}(0,{\bm{I}})$. The superscript “*” indicates either “$w$” (for the preferred data) or “$l$” (for the less preferred data). The expectation is taken over samples $\{{\bm{x}}_{0}^{w},{\bm{x}}_{0}^{l}\}\sim\mathcal{D}$ and the noise schedule $t$. In Rectified Flow, we relate the noise vector ${\bm{\epsilon}}^{*}$ to a velocity field ${\bm{v}}^{*}$. Specifically, Lemma A.1 in Appendix A shows that
@@
where ${\bm{\epsilon}}_{\text{pred}}$ and ${\bm{v}}_{\text{pred}}$ refer to predictions either from the model $p_{\theta}$ or the reference model $p_{\text{ref}}$. By substituting Eq. 9 into Eq. 4.1, we obtain the final Flow-DPO loss $\mathcal{L}_{\text{FD}}(\theta)$:
@@
$\quad-\bigl{(}\|{\bm{v}}^{l}-{\bm{v}}_{\theta}({\bm{x}}_{t}^{l},t)\|^{2}-\|{\bm{v}}^{l}-{\bm{v}}_{\text{ref}}({\bm{x}}_{t}^{l},t)\|^{2}\bigr{)}\Big{)}\Bigg{)}\Bigg{]},$ (10)
where $\beta_{t}=\beta\,(1-t)^{2}$ and the expectation is taken over samples $\{{\bm{x}}_{0}^{w},{\bm{x}}_{0}^{l}\}\sim\mathcal{D}$ and the noise schedule $t$. Intuitively, minimizing $\mathcal{L}_{\text{FD}}(\theta)$ guides the predicted velocity field ${\bm{v}}_{\theta}$ closer to the target velocity ${\bm{v}}^{w}$ of the “preferred” data, while pushing it away from ${\bm{v}}^{l}$ (the “less preferred” data). The strength of this preference signal depends on the differences between the predicted errors and the corresponding reference errors, $\|{\bm{v}}^{w}-{\bm{v}}_{\text{ref}}({\bm{x}}_{t}^{w},t)\|^{2}$ and $\|{\bm{v}}^{l}-{\bm{v}}_{\text{ref}}({\bm{x}}_{t}^{l},t)\|^{2}$. We provide a pseudo-code of Flow-DPO in Appendix E.
@@
The parameter $\beta_{t}$ governs the strength of the KL divergence constraint, determining the extent to which the model can diverge from the base reference model (Rafailov et al., 2024; Wallace et al., 2024). Note that from our derivation, $\beta_{t}=\beta(1-t)^{2}$. As the timestep $t$ approaches 1, $\beta_{t}$ approaches zero, resulting in a relatively weak KL constraint. Conversely, as $t$ approaches 0, $\beta_{t}$ converges to the original $\beta$, enforcing a stronger KL constraint. This scheduling strategy causes the model to prioritize alignment at higher noise levels. Our experiments reveal that this degrades performance, while using a constant $\beta_{t}$ provides better results. Similar phenomena have been observed in DDPM’s (Ho et al., 2020) standard diffusion training objective, where discarding the weighting in denoising score matching leads to improved sample quality, as discussed in Section 3.4 of their paper. We provide a more detailed discussion of this in Section 5.2.
@@ 4.2 Flow-RWR
Drawing inspiration from the application of Reward-weighted Regression (RWR) (Peters & Schaal, 2007) in diffusion models (Lee et al., 2023; Furuta et al., 2024), we propose a counterpart for flow-based models based on expectation-maximization. Starting from the general KL-regularized reward-maximization problem in Eq. 4.1, prior work (Rafailov et al., 2024; Peters & Schaal, 2007) shows that its optimal closed-form solution can be written as
$p_{\theta}(\mathbf{x}_{0}\mid\mathbf{y})=\frac{1}{Z(\mathbf{y})}\,p_{\text{ref}}(\mathbf{x}_{0}\mid\mathbf{y})\,\exp\biggl{(}\frac{1}{\beta}\,r(\mathbf{x}_{0},\mathbf{y})\biggr{)},$ (11)
where $Z(\mathbf{y})=\sum_{\mathbf{x}_{0}}p_{\text{ref}}(\mathbf{x}_{0}\mid\mathbf{y})\exp\bigl{(}\tfrac{1}{\beta}r(\mathbf{x}_{0},\mathbf{y})\bigr{)}$ is the partition function. Following (Furuta et al., 2024), we can obtains the RWR loss by M-step:
$\mathcal{L}_{\text{RWR}}(\theta)=-\mathbb{E}_{{\bm{y}}\sim\mathcal{D}_{c}}\bigl{[}\frac{1}{Z({\bm{y}})}\exp(\frac{r({\bm{x}}_{0},{\bm{y}})}{\beta})\log p_{\theta}({\bm{x}}_{0}\mid{\bm{y}})\bigr{]}.$
In practice, we omit the intractable normalization $Z(\mathbf{y})$ and $\beta$, and we employ Eq. 37 in Wallace et al. (2024) as a simplified upper bound of the negative log-likelihood. This yields the approximate form
@@
Substituting Eq. 9 into Eq. 12 then provides the final RWR objective for velocity prediction:
@@
where $\mathbf{v}$ is the ground-truth velocity, and $\mathbf{v}_{\theta}(\mathbf{x}_{t},t,\mathbf{y})$ represents the predicted velocity under parameters $\theta$. We omit $(1-t)^{2}$ as in Sec. 4.1. By applying RWR in this way, our flow-based model learns to generate velocity fields that emphasize high-reward samples, paralleling how diffusion-based models align samples via reward-weighted likelihood maximization.
@@ 4.3 Noisy Reward Guidance
Recall that Eq. 4.1 has a closed-form solution as presented in Eq. 11. This expression indicates that the goal of RLHF is to transform the original distribution $p_{\text{ref}}({\bm{x}}_{0}\mid{\bm{y}})$ into the new target distribution $p_{\theta}({\bm{x}}_{0}\mid{\bm{y}})$. Since the constants $\beta$ and $w$ can be absorbed into $r({\bm{x}}_{0},{\bm{y}})$, Eq. 11 becomes
@@
where $w\in\mathbb{R}$ controls the strength of the reward guidance. Notably, this formulation aligns with the concept of classifier guidance (Dhariwal & Nichol, 2021; Song et al., 2020), so we refer to it as reward guidance.
As we proved in Appendix A.2, for Rectified Flow, if we set the marginal velocity field to
@@
the marginal distribution is modified from $p_{t}({\bm{x}}_{t}\mid{\bm{y}})$ to
$\tilde{p}_{t}({\bm{x}}_{t}\mid{\bm{y}})\;\propto\;p_{t}({\bm{x}}_{t}\mid{\bm{y}})\,[\exp(r({\bm{x}}_{t},{\bm{y}}))]^{w}.$
When $t=0$, the resulting samples are drawn from Eq. 14. We provide a pseudo-code in Appendix E. To producing meaningful reward values for noised videos ${\bm{x}}_{t}$, we train a time-dependent reward function. We use the same dataset $D=\{{\bm{y}},{\bm{x}}^{w},{\bm{x}}^{l}\}$ described in Section 3.2, where ${\bm{x}}^{w}$ and ${\bm{x}}^{l}$ represent two videos with different preference levels. In order to train our reward function, we assume that applying the same noise to both videos preserves their preference relationship. Formally, for each pair $({\bm{x}}^{w},{\bm{x}}^{l})$, we define
${\bm{x}}_{t}^{w}=(1-t)\,{\bm{x}}_{0}+t\,{\bm{x}}_{1}^{w},$
${\bm{x}}_{t}^{l}=(1-t)\,{\bm{x}}_{0}+t\,{\bm{x}}_{1}^{l},\text{with}\,{\bm{x}}_{t}^{w}\succ{\bm{x}}_{t}^{l}.$ (16)
Here, ${\bm{x}}_{0}$ is the original video, while ${\bm{x}}_{1}^{w}$ and ${\bm{x}}_{1}^{l}$ are the noised versions corresponding to the more preferred and less preferred videos, respectively. We then adopt the BT model to learn the reward function from these noised videos. Specifically, we minimize the following loss:
$\mathcal{L}_{BT}(\theta)=-\mathbb{E}\bigl{[}\log\bigl{(}\sigma\bigl{(}r_{\theta}({\bm{x}}_{t}^{w},t)-r_{\theta}({\bm{x}}_{t}^{l},t)\bigr{)}\bigr{)}\bigr{]},$ (17)
Table 2: Preference accuracy on GenAI-Bench and VideoGen-Eval. w/ Ties indicates that accuracy is calculated with ties included (Deutsch et al., 2023), and w/o Ties excludes tied pairs when calculating accuracy. * denotes that for LiFT, ties prediction are randomly converted to chosen/rejected with a 0.5 probability due to a large number of ties produced by the model. Bold: best performance.

### v2

@@ 3.2 Reward Modeling
where $c$ denotes preference choice, $\theta>1$ controls the tendency toward ties, with a larger $\theta$ increasing tie probability. We empirically set $\theta=5.0$ and train the BTT model by minimizing negative log-likelihood:
$\mathcal{L}_{\text{BTT}}=-\mathbb{E}_{({\bm{y}},{\bm{x}}_{0}^{A},{\bm{x}}_{0}^{B})\sim\mathcal{D}}\Big[\sum_{c\in\{\succ,\prec,=\}}\mathbbm{1}(c)\log P_{\theta}(c|{\bm{y}},{\bm{x}}_{0}^{A},{\bm{x}}_{0}^{B})\Big]$ (2)
We train BT and BTT reward models under identical settings and visualize $\Delta r=r({\bm{x}}_{0}^{A},{\bm{y}})-r({\bm{x}}_{0}^{B},{\bm{y}})$ on the validation set (Fig. 3). Although the BT model handles chosen/rejected pairs well, it struggles to handle ties—often assigning sizeable $\Delta r$ to many tie pairs, conflating them with clear preferences. By contrast, BTT learns a flexible decision boundary that clusters ties near zero while preserving large margins for decisive wins and losses, yielding more reliable feedback for downstream RLHF.
Token Positioning.
A common approach in LLM / MLLM-based reward modeling [52, 69, 19] attaches a linear projection head to the final token to predict multi-dimensional scores. This method forcing prompt-independent and prompt-dependent cues into one vector, causing context leakage: the same video can receive different visual-quality scores when paired with different prompts. We eliminate this entanglement with a simple token-positioning strategy. As shown in Fig.1(b), two context-agnostic tokens, [VQ], [MQ], are inserted immediately after the video and before the prompt, so they can attend only to visual content. A context-aware token, [TA], is placed after the full prompt, allowing it to attend to both the video and the text. The final-layer embeddings of these tokens are then mapped to dimension-specific scores via a shared linear layer. This design removes context leakage, stabilizes visual and motion assessments, and maintains parameter efficiency. The full input template is provided in Appendix K.
@@ 4 Video Alignment
With a high-fidelity reward model established, we next introduce three alignment methods for flow-based generation models under a unified RL objective: two training-time algorithms—Flow-DPO and Flow-RWR, and one inference-time guidance technique, Flow-NRG (Fig. 1 (c)).
4.1 Preliminaries
Rectified Flow.
Let ${\bm{x}}_{0}\sim q({\bm{x}}_{0})$ denote a data sample from the real data distribution, and ${\bm{x}}_{1}\sim p({\bm{x}}_{1})$ denote a noise sample, where ${\bm{x}}_{0},{\bm{x}}_{1}\in\mathbb{R}^{d}$. Recent advanced image [15] and video generation models [57, 32] adopt the Rectified Flow [48], which defines the “noisy” data ${\bm{x}}_{t}$ as ${\bm{x}}_{t}=(1-t)\,{\bm{x}}_{0}\;+\;t\,{\bm{x}}_{1},$ where $t\in[0,1]$. Then we can train a transformer model to regress the velocity field ${\bm{v}}_{\theta}({\bm{x}}_{t},t)$ by minimizing the Flow Matching objective [45, 48]:
@@ 4.1 Preliminaries
RLHF.
The goal of RLHF is to learn a conditional distribution $p_{\theta}({\bm{x}}_{0}\mid{\bm{y}})$ that maximizes the reward $r({\bm{x}}_{0},{\bm{y}})$ while controls the KL-divergence from the reference model $p_{\text{ref}}$ via a coefficient $\beta$:
$\max_{p_{\theta}}\ \mathbb{E}_{{\bm{y}}\sim\mathcal{D}_{c},{\bm{x}}_{0}\sim p_{\theta}({\bm{x}}_{0}\mid{\bm{y}})}\left[r({\bm{x}}_{0},{\bm{y}})\right]-\beta\,\mathbb{D}_{\text{KL}}\left[p_{\theta}({\bm{x}}_{0}\mid{\bm{y}})\,\|\,p_{\text{ref}}({\bm{x}}_{0}\mid{\bm{y}})\right].$ (3)
@@ 4.2 Flow-DPO
Consider a training set $\mathcal{D}=\{{\bm{y}},{\bm{x}}_{0}^{w},{\bm{x}}_{0}^{l}\}$, where $y$ is the prompt and human annotations prefer generated video ${\bm{x}}_{0}^{w}$ to ${\bm{x}}_{0}^{l}$ (i.e., ${\bm{x}}_{0}^{w}\succ{\bm{x}}_{0}^{l}$). Direct Preference Optimization (DPO) [61] aligns models with human preferences by analytically solving the RLHF objective (Eq. 3) and optimizing the policy via supervised learning. Extending this idea to diffusion models, Diffusion-DPO [71] derives a DPO-style loss under the diffusion paradigm. The resulting objective $\mathcal{L}_{\text{DD}}(\theta)$ is formulated as:
$\mathbb{E}\Bigg[\log\sigma\Bigg(-\frac{\beta}{2}\Big(\|{\bm{\epsilon}}^{w}-{\bm{\epsilon}}_{\theta}({\bm{x}}_{t}^{w},t)\|^{2}-\|{\bm{\epsilon}}^{w}-{\bm{\epsilon}}_{\text{ref}}({\bm{x}}_{t}^{w},t)\|^{2}-\bigl(\|{\bm{\epsilon}}^{l}-{\bm{\epsilon}}_{\theta}({\bm{x}}_{t}^{l},t)\|^{2}-\|{\bm{\epsilon}}^{l}-{\bm{\epsilon}}_{\text{ref}}({\bm{x}}_{t}^{l},t)\|^{2}\bigr)\Big)\Bigg)\Bigg]$ (4)
The expectation is taken over samples $\{{\bm{x}}_{0}^{w},{\bm{x}}_{0}^{l}\}\sim\mathcal{D}$ and the noise schedule $t$. In Rectified Flow, we relate the noise vector $\bm{\epsilon}$$^{*}$ to a velocity field $\bm{v}$$^{*}$. Specifically, Lemma B.1 in Appendix B shows that
@@
By substituting Eq. 5 into Eq. 4, we obtain the final Flow-DPO loss $\mathcal{L}_{\text{FD}}(\theta)$:
@@
$\quad-\bigl(\|{\bm{v}}^{l}-{\bm{v}}_{\theta}({\bm{x}}_{t}^{l},t)\|^{2}-\|{\bm{v}}^{l}-{\bm{v}}_{\text{ref}}({\bm{x}}_{t}^{l},t)\|^{2}\bigr)\Big)\Bigg)\Bigg]$ (6)
where $\beta_{t}=\beta\,(1-t)^{2}$ . Intuitively, minimizing $\mathcal{L}_{\text{FD}}(\theta)$ guides the predicted velocity field ${\bm{v}}_{\theta}$ closer to the target velocity ${\bm{v}}^{w}$ of the “preferred” data, while pushing it away from ${\bm{v}}^{l}$ (the “less preferred” data). The strength of this preference signal depends on the differences between the predicted errors and the corresponding reference errors. We provide Flow-DPO pseudo-code in Appendix C.
@@
The KL coefficient $\beta_{t}$ controls how far the learned policy is allowed to deviate from the reference model [61, 71]. A direct derivation yields the schedule $\beta_{t}=\beta(1-t)^{2}$. The penalty $\beta_{t}$ vanishes as $t$ approaches 1 and reaches $\beta$ at $t=0$. This scheduling strategy causes the model to prioritize alignment at higher noise levels. Unlike in diffusion models, our experiments reveal that this schedule degrades alignment performance in rectified flow, leading to reward hacking and visual artifacts. Inspired by a similar observation in DDPM’s [22] training objective, where discarding the weighting in denoising score matching improves sample quality, we instead adopt a constant $\beta$. This adjustment leads to more stable training and improved alignment across all reward dimensions. We provide a more detailed discussion of this in Section 5.2.
@@ 4.3 Flow-RWR
Drawing inspiration from the application of Reward-weighted Regression (RWR) [55] in diffusion models [36, 17], we propose a counterpart for flow-based models based on expectation-maximization [10]. Starting from the general KL-regularized reward-maximization problem in Eq. 3, prior work [61] shows that its optimal closed-form solution can be written as:
$p_{\theta}({\bm{x}}_{0}\mid{\bm{y}})=\frac{1}{Z({\bm{y}})}\,p_{\text{ref}}({\bm{x}}_{0}\mid{\bm{y}})\,\exp\biggl(\frac{1}{\beta}\,r({\bm{x}}_{0},{\bm{y}})\biggr),$ (7)
where $Z({\bm{y}})=\sum_{{\bm{x}}_{0}}p_{\text{ref}}({\bm{x}}_{0}\mid{\bm{y}})\exp\bigl(\tfrac{1}{\beta}r({\bm{x}}_{0},{\bm{y}})\bigr)$ is the partition function. Following [17], we can obtains the RWR loss:
@@
For rectified-flow models, we formulate this as a reward-weighted velocity regression:
@@
As in Flow-DPO, we omit the $(1-t)^{2}$ factor for better performance.
@@ 4.4 Noisy Reward Guidance
Recall that the KL-regularized RL objective (Eq. 3) admits a closed-form solution(Eq. 7), which transform the original distribution $p_{\text{ref}}({\bm{x}}_{0}\mid{\bm{y}})$ into the new target distribution $p_{\theta}({\bm{x}}_{0}\mid{\bm{y}})$. Since the constants $\beta$ and $w$ can be absorbed into $r({\bm{x}}_{0},{\bm{y}})$, the closed-form solution becomes:
@@
where $w\in\mathbb{R}$ controls the strength of the reward guidance. For rectified flow, as we proved in Appendix B.2, this reweighting can be achieved by shifting the velocity field:
@@
This modification of the marginal velocity field alters the sampling distribution to match the target form in Eq. equation 10. Since this formulation is structurally similar to classifier guidance [12, 67], we refer to it as reward guidance. Pseudo-code is provided in Appendix C.
Efficient Reward on Noisy Latents.
Computing $\nabla r$ in pixel space requires back-propagating through the full VAE decoder, which is computationally expensive. To address this, we propose training a lightweight, time-dependent reward model $r_{\theta}(\cdot,t)$ directly in latent space. For each preference pair $({\bm{x}}^{w},{\bm{x}}^{l})$, we apply identical noise to both videos, assuming their relative preference remains unchanged. We then adopt the Bradley–Terry loss to learn the reward function from these noised videos. Leveraging the fact that modern VDMs are already well trained on noisy latents, we can reuse a few early layers from the pretrained backbone to construct the reward model, avoiding the need for heavy retraining. We apply Eq. equation 11 at each inference step (except at $t=1$), enabling efficient inference-time alignment in latent space.
5 Experiments
Table 2: Preference accuracy on GenAI-Bench and VideoGen-RewardBench. For ties-excluded accuracy, we calculate accuracy using only the data labeled as “A wins” or “B wins”. For ties-included accuracy, we use the algorithm from Deutsch et al. [11], which tests various tie thresholds and selects the one that maximizes three-class accuracy. Bold: best performance.